# Early Warning Signals for Bifurcations Embedded in High Dimensions

### v2

@@ Introduction
Figure 6: Comparison between the neural model’s EWS estimation for meridional ($v$) vs. zonal ($u$) flow. The thick lines represent median output over all years 1940-2022, and the shaded regions denote the 25th and 75th percentiles over the same set. The approximate date of the monsoon shift is denoted by the dashed vertical line
The strength of these results for both the $u$ and $v$ components of neutral wind merits a brief discussion regarding variable selection in these spatially-informed EWS methods. The results presented in this paper have focused primarily on inferring a bifurcation embedding across the degrees of freedom of a single continuum variable, but the same approach can be applied across multiple dynamic variables being measured simultaneously. Interpretation of multivariate results introduces some additional subtleties, however. We have thus far assumed that positive identification of early warning indicators along a given axis implies its non-orthogonality to the embedded bifurcation manifold, but this is not always the case. Consider a collection of system variables $\mathbf{x}$ with some bifurcation dynamics given by $\dot{\mathbf{x}}=\mathbf{f}\left(\mathbf{x}\right)$, and an ancillary variable space $\mathbf{y}$ evolving as $\dot{\mathbf{y}}=\mathbf{h}\left(\mathbf{x},\mathbf{y}\right)$. The dependence of $\mathbf{y}$ on $\mathbf{x}$ means that critical slowing down phenomena (e.g.) in $\mathbf{x}$ as it approaches its critical transition may propagate to $\mathbf{y}$, even though it is entirely exogeneous to the bifurcation system. The problem posed by these “downstream” variables is typically negligible when comparing degrees of freedom of one continuum variable, since coupling is generally symmetric (i.e. the extent of influence of dynamics at location A on those at location B is expected to be equal to that of B on A). This assumption of symmetry may not hold when comparing two different variables, however, which can complicate the interpretation of EWS results. The criteria for the presence of EWS along a given direction may be better understood using the observability framework of control theory, as suggested in [37]. For the sake of simplicity, we restrict our analysis in this work to variables that can reasonably be assumed to couple approximately symmetrically (e.g. the vertical and horizontal flow fields).
@@
Figure 7: Comparison between the neural model’s EWS estimation for the two meridional wind variables. The thick lines represent median output over all years 1940-2022, and the shaded regions denote the 25th and 75th percentiles over the same set. The approximate date of the monsoon shift is denoted by the dashed vertical line
@@ Conclusions
The practical potential of this approach is particularly salient in the study of Earth systems, where characterizing potential tipping points has become an increasingly urgent topic of research as the effects of climate change continue to intensify. The West African monsoon shift made for a useful test case in this study because its mechanism is already somewhat well understood and has been observed repeatedly, allowing for validation of the results obtained. We hope, however, that the same methodology can prove useful in identifying as-yet unobserved tipping phenomena and in helping to refine predictive models of how these events might unfold, both in the climate and other complex nonlinear systems.
Acknowledgments
Data for the turbulent channel flow system is made available by the Johns Hopkins Turbulence Database at http://turbulence.pha.jhu.edu. All data used for the West African Monsoon Shift analysis comes from the ERA5 global reanalysis [35] provided by the Copernicus Climate Change Service (C3S) Climate Data Store (CDS) at https://doi.org/10.24381/cds.143582cf [36]. Neither the European Commission nor ECMWF is responsible for any use that may be made of the Copernicus information or data it contains.
Funding
The research was supported by NSERC Discovery Grants to MA (5006032-2016) and CTB (5013291-2019), and a DARPA Artificial Intelligence Exploration Opportunity Grant to MA and CTB (PA-21-04-02-ACTM-FP-012).
Appendix A EWS Detection Model
@@
The neural EWS detection model used in this work is inspired by that of Bury et. al. (2021) [14]. Taking a one-dimensional time series as input, the network consists of two convolutional layers of 20 units each (kernel width 8) followed by an LSTM module with 20 memory cells. Between the CNN and LSTM layers there is a 10% dropout layer and a 1D max pooling layer of width 2. Finally, a dense softmax layer is used to produce probability-like outputs for each of the target classes. A diagram of the model’s architecture is shown in Fig. 8.
@@
Figure 8: Schematic representation of the neural network used for EWS detection. The dimensions labeled ”None” refer to the user-specified batch size
@@
Appendix B Mean-Field Model for Vortex Shedding System
The mean-field solution introduced in [18] identifies three spatial modes ($u_{x},u_{y},u_{w}$, illustrated in Fig. 1) relevant to the bifurcation dynamics. Projecting the system’s high-dimensional evolution into the 3D subspace of these modes yields dynamics which are, to good approximation, described by the following interactions:
$\begin{split}\dot{w}&=-\lambda\left(w-x^{2}-y^{2}\right)\\ \dot{x}&=\mu x-\eta y+Axw\\ \dot{y}&=\eta x+\mu y+Ayw\\ \end{split}$ (3)
For large values of $\lambda$ this system decomposes into two distinct timescales: the first equation leads $w$ to quickly converge to $w=x^{2}+y^{2}$, and the remaining slow dynamics unfold on the resulting two-dimensional manifold. Substituting $w=x^{2}+y^{2}$ into the other equations and introducing the variable $z=x+iy$, the slow dynamics reduce to
$\begin{split}\dot{z}&=z\left((\mu+i\eta)+A|z|^{2}\right)\\ \end{split}$ (4)
which is the normal form for a Hopf bifurcation.

### v3

@@ Results
Figure 6: Comparison between the neural model’s EWS estimation for meridional ($v$) vs. zonal ($u$) flow. The thick lines represent median output over all years 1940-2022, and the shaded regions denote the 25th and 75th percentiles over the same set. Both signals increase significantly in the weeks leading up to the approximate date of the monsoon shift, which is denoted by the dashed vertical line.
The strength of these results for both the $u$ and $v$ components of neutral wind merits a brief discussion regarding variable selection in these spatially-informed EWS methods. The results presented in this paper have focused primarily on inferring a bifurcation embedding across the degrees of freedom of a single continuum variable, but the same approach can be applied across multiple dynamic variables being measured simultaneously. Interpretation of multivariate results introduces some additional subtleties, however. We have thus far assumed that positive identification of early warning indicators along a given axis implies its non-orthogonality to the embedded bifurcation manifold, but this is not always the case. Consider a collection of system variables $\mathbf{x}$ with some bifurcation dynamics given by $\dot{\mathbf{x}}=\mathbf{f}\left(\mathbf{x}\right)$, and an ancillary variable space $\mathbf{y}$ evolving as $\dot{\mathbf{y}}=\mathbf{h}\left(\mathbf{x},\mathbf{y}\right)$. The dependence of $\mathbf{y}$ on $\mathbf{x}$ means that critical slowing down phenomena (e.g.) in $\mathbf{x}$ as it approaches its critical transition may propagate to $\mathbf{y}$, even though it is entirely exogeneous to the bifurcation system. The problem posed by these “downstream" variables is typically negligible when comparing degrees of freedom of one continuum variable, since coupling is generally symmetric (i.e. the extent of influence of dynamics at location A on those at location B is expected to be equal to that of B on A). This assumption of symmetry may not hold when comparing two different variables, however, which can complicate the interpretation of EWS results. The criteria for the presence of EWS along a given direction may be better understood using the observability framework of control theory, as suggested in [34]. For the sake of simplicity, we restrict our analysis in this work to variables that can reasonably be assumed to couple approximately symmetrically (e.g. the vertical and horizontal flow fields).
@@
Figure 7: Comparison between the neural model’s EWS estimation for the two meridional wind variables. The thick lines represent median output over all years 1940-2022, and the shaded regions denote the 25th and 75th percentiles over the same set. The approximate date of the monsoon shift is denoted by the dashed vertical line. Although both signals exhibit a positive trend in the lead-up to the monsoon shift, the increase is much more prominent in the case of the 10m neutral wind.
@@ Conclusions
Although the methods presented are generically applicable to any high dimensional systems undergoing local bifurcations, their efficacy is constrained by the shortcomings of the chosen EWS detection technique—all EWS algorithms have different strengths and weaknesses with respect to accuracy, generalizability, and interpretability. Additionally, the issue of nonuniform detection sensitivity due to spatially heterogeneous perturbations may in some cases pose further difficulty to the data-driven inference process. In spite of these limitations, the demonstrated utility of these methods on both a simple synthetic system and an empirically observed climate phenomenon suggests that they have a practical role to play in the analysis of real-world bifurcation events.
@@
The pragmatic potential of this approach is particularly salient in the study of Earth systems, where characterizing potential tipping points has become an increasingly urgent topic of research as the effects of climate change continue to intensify. The West African monsoon shift made for a suitable test case in this study because its mechanism is already somewhat well understood and has been observed repeatedly, allowing for validation of the results obtained. We hope, however, that the same methodology can prove useful in identifying as-yet unobserved tipping phenomena and in helping to refine predictive models of how these events might unfold, both in the climate and other complex nonlinear systems.
Methods
EWS Detection Methods
Results presented in this work are based on the application of an EWS detection model to time series data. A large variety of methodologies have been proposed for measuring signatures of oncoming critical transitions, diverse in both the types of transitions they are intended to forecast (local bifurcations, nonlocal bifurcations, phase transitions, etc.) [7, 35, 36] and in the analytical means by which they do so (temporal or spatial computation of autocorrelation, moments of distribution, spectral methods, etc.) [8, 37]. Additionally, recent work has shown the potential of machine learning models to improve on these traditional statistical approaches by improving on accuracy and offering greater insight into the nature of the oncoming transition [11, 38, 16]. In this paper we employ a neural model of our own construction, based primarily on the design of Bury et. al. (2021) [11]. The classifier takes as input a univariate time series of arbitrary length, and outputs a vector of probabilities assigned to each of the target classes: null (no upcoming bifurcation) or any of four types of local bifurcation (fold, Hopf, transcritical, and pitchfork).
We chose a deep learning approach for its classification accuracy and flexibility, but these methods are not without drawbacks. Neural networks are well known to be fairly opaque in their execution, sacrificing interpretability for performance. Furthermore, they tend to overstate their confidence in results for classification tasks, i.e. outputting class probabilities very close to 0% or 100% even when their demonstrated accuracy on validation tasks is considerably lower (though this can be mitigated somewhat through calibration methods such as temperature scaling) [39]. We emphasize that none of the analysis in this paper is specific to any one EWS detection method. All results we present are produced by our neural classifier model, but the same methodologies could be equally effectively carried out using any other approach (modulo any potential shortcomings of the chosen technique).
Neural EWS Model
@@ Methods
The neural EWS detection model used in this work is based on that of Bury et. al. (2021) [11], code for which is available at https://github.com/ThomasMBury/deep-early-warnings-pnas. Taking a one-dimensional time series as input, the network consists of two convolutional layers of 20 units each (kernel width 8) followed by an LSTM module with 20 memory cells. Between the CNN and LSTM layers there is a 10% dropout layer and a 1D max pooling layer of width 2. Finally, a dense softmax layer is used to produce probability-like outputs for each of the target classes. A diagram of the model’s architecture is shown in Fig. 8.
@@
Figure 8: Schematic representation of the neural network used for EWS detection. The dimensions labeled "None" refer to the user-specified batch size.